If α , β are roots of the equation 9x 2 +6x+1=0 then find the equation whose roots are reciprocal to the given roots.

 If  α , β are roots of the equation 9x2+6x+1=0 then find the equation whose roots are reciprocal to the given roots.

Grade:12th pass

1 Answers

Ass satwik
5 years ago
f(x)= 9x^2+6x+1 
Now we have to find a equation whose root is y where y=(1/x) 
So x=(1/y)
Substituting in f(x) we get 
f(1/y)= 9(1/y)^2+6(1/y)+1
Simplifing will give y^2+6y+9 
So, x^2+6x+9=0 is equation with reciprocal roots of 9x^2+6x+1=0.